OpenStudy (anonymous):

What is the approximate circumference of a circle with a diameter of 56 cm? (Use π = 3.14.) A. 87.92 cm B. 175.84 cm C. 351.68 cm D. 10.82 cm

OpenStudy (anonymous):

C =2(pi)(r)

OpenStudy (anonymous):

divide diameter by 2 to get radius

OpenStudy (anonymous):

so 56/2 =28

OpenStudy (anonymous):

r =28

OpenStudy (anonymous):

C = 2(3.14)(28)

OpenStudy (anonymous):

C = (6.28)(28)

OpenStudy (anonymous):

C =175.84

OpenStudy (anonymous):

So, the answer is B